On December 31, 2020, Blossom Company acquired a computer from Plato Corporation by issuing a $650,000 zero-interest-bearing note, payable in full on December 31, 2024. Blossom Company’s credit rating permits it to borrow funds from its several lines of credit at 12%. The computer is expected to have a 5-year life and a $76,000 salvage value.

Prepare the journal entry for the purchase on December 31, 2020. (Round present value factor calculations to 5 decimal places, e.g. 1.25124 and the final answers to 0 decimal places e.g. 58,971. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually.)

Date

Account Titles and Explanation

Debit

Credit

December 31, 2020

Prepare any necessary adjusting entries relative to depreciation (use straight-line) and amortization (use effective-interest method) on December 31, 2021. (Round answers to 0 decimal places, e.g. 38,548. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually.)

Date

Account Titles and Explanation

Debit

Credit

December 31, 2021

(To record the depreciation.)

December 31, 2021

(To amortize the discount.)

Schedule of Note Discount Amortization


Date

Debit, Interest Expense Credit,
Discount on Notes Payable

Carrying Amount
of Note

12/31/20 $

$

12/31/21

12/31/22

12/31/23

12/31/24

Prepare any necessary adjusting entries relative to depreciation and amortization on December 31, 2022. (Round answers to 0 decimal places, e.g. 38,548. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually.)

Date

Account Titles and Explanation

Debit

Credit

December 31, 2022

(To record the depreciation.)

December 31, 2022

(To amortize the discount.)
Answer #1

Working as follows:

Amortization Schedule
Date Interest Expense Carrying Amount of Note
12/31/2020 $413,087
12/31/2021 $413087 × 12% = $49570 $462,657
12/31/2022 $462,657 × 12% = $55,519 $518,176
12/31/2023 $518,176 × 12% = $62,181 $580,357
12/31/2024 $580,357 × 12% = $69,643 $650,000

_________________________________________________________________________

Prepare Journal entries for purchase as follows:

Date Account Titles Debit Credit
Dec. 31, 2020 Computer Equipment [$650,000 × 0.635518 $413,087
Discount on notes payable $236,913
Note payable $650,000

Using PVF(12%, 4) = 0.635518

______________________________________________________________________________

Prepare any necessary adjusting entries relative to depreciation and amortization:

Date Account Titles Debit Credit
Dec. 31, 2021 Depreciation Expense $67,417.40
Accumulated Depreciation- Computer equip. $67,417.40
($413,087-$76,000)/5 years
Dec. 31, 2021 Interest Expense $49,570
Discount on Notes payable $49,570

___________________________________________________________________________

Prepare adjusting entries relative to depreciation and amortization on December 31, 2022 as follows:

Date Account Titles Debit Credit
Dec. 31, 2022 Depreciation Expense $67,417.40
Accumulated Depreciation- Computer equip. $67,417.40
($413,087-$76,000)/5 years
Dec. 31, 2022 Interest Expense $55,519
Discount on Notes payable $55,519